OpenStudy (anonymous):

jack wants to know how many families in his neighborhood of 60 homes would help organize a neighborhood fund raising party. He put all the addresses in a bag and drew a random sample of 30 addresses. He then asked those families if they would help organize the fund raising party. He found that 12% of the families would help organize the party. He claims that 12% of the neighborhood families would be expected to help organize the party. is this a valid inference? A.No,this is not valid because he asked only 30 families B.No,this is not valid because he did not take a random sample the the neig

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I think it is A, since 12% of 30 families, of 60 homes, is: \[\frac{{12}}{{100}} \times \frac{{50}}{{100}} = \frac{{600}}{{10000}} = \frac{6}{{100}} = 6\% \] of the whole sample

OpenStudy (anonymous):

okay thx ^.^

OpenStudy (michele_laino):

thanks!